OpenStudy (anonymous):

3sqrt2x^3 + 4xsqrt8x

OpenStudy (lgbasallote):

hmm..first ill ask you...where are you having difficulties with in there?

OpenStudy (anonymous):

Understanding the whole concept. The teacher just gives the answer and doesnt explain the process

OpenStudy (lgbasallote):

hmmm first thing you might wanna do is factor out the values inside of the square roots

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

\[3\sqrt{2x^3} = 3\sqrt{2x \times x^2} = 3x \sqrt{2x}\] \[4x \sqrt{8x} = 4x \times 2 \sqrt{2x}\] so the problem is now \[3x \sqrt{2x} + 8x \sqrt{2x} = 11x \sqrt{2x}\]